# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 3.3.2. Multi-Head Structured Attention
(6) $q_{i}=\sigma(\mathbf{W}^{\mathsf{T}}\cdot(\hat{x}_{i};\hat{x}_{s})+b),$
@@ 3.3.3. Similarity-based Representation Aggregation
In the detection phase, we sample $n$ other message nodes for each message node $m_{i}$. Consequently, for each specific message $m_{i}$, we obtain a set of $n$ candidate representations $X_{i}=[\hat{x}_{i}^{0},\hat{x}_{i}^{1},\cdots,\hat{x}_{i}^{n-1}]$, along with the corresponding set of similarity $Q_{i}=[q_{i}^{0},q_{i}^{1},\cdots,q_{i}^{n-1}]$. Subsequently, select the indices from $Q_{i}$ that satisfy the conditions and average the candidate representations corresponding to these indices to obtain the final representation of $m_{i}$. We formalize this process as follows:
(7) $I_{i}=\{j\mid q_{i}^{j}>\alpha,j\in\{0,1,\ldots,n-1\}\},\quad\overline{x}_{i}=\frac{1}{|I_{i}|}\sum_{j\in I_{i}}\hat{x}_{i}^{j}.$
Here, $I_{i}$ represents the set of indices obtained by filtering the set $Q_{i}$, $\alpha$ is the similarity retention threshold and $\overline{x}_{i}$ denotes the final representation of message $m_{i}$. Particularly, we choose the candidate representation corresponding to the highest similarity index in $Q_{i}$ as the final representation for those messages in the similarity set $Q_{i}$ with only one or no indices meeting the condition.
@@ 3.4.2. Clustering Constraint
Input: A data block: $D$, batch size: $B$, whether message block is a training dataset: Train.
Output: Candidate representation sets of messages: $X_{1},X_{2},X_{3},\cdots$, and their corresponding similarity sets: $Q_{1},Q_{2},Q_{3},\cdots$.
1 $batches=[D^{1},D^{2},D^{3},\cdots]_{B}$; // Split dataset into batches of size $B$
2 for batch in batches do
3       Get representations of messages and sampled messages via Eq. 5;
4       Get the similarities of message pairs via Eq. 6;
5       if Train == True then
@@ 9             Back propagate to update model parameters;
11             Update all sets of message candidate representations: $X_{1},X_{2},\cdots,X_{i}$;
12             Update all similarity sets: $Q_{1},Q_{2},\cdots,Q_{i}$;
13             Execute Algorithm 4;
@@
Algorithm 3 Multi-relational Prompt-based Pairwise Message Learning Mechanism.
@@ 3.5. Proposed $\mathrm{RPLM}_{SED}$
As depicted in Algorithm 1, the life cycle of $\mathrm{RPLM}_{SED}$ is divided into three stages: the initial training phase, the detection phase, and the maintenance phase. During the initial training stage, we train the model using the initial message block (Algorithm 1, lines 4, 11-16). Subsequently, the model performs event detection on message blocks within the window (Algorithm 1, lines 18-23). To retain existing knowledge while adapting to new data, we perform a weighted update of the parameters in the PLMs within the model before the maintenance phase (Algorithm 1, line 6-8), then maintain the model by utilizing the message blocks within the window (Algorithm 1, lines 9, 11-16). Once the model maintenance is completed, it will proceed to detect events in message blocks within the next window. Algorithm 2 details the pairwise message modeling strategy, including the message node sampling and multiple relation mapping (Algorithm 2, lines 7 to 19). In Algorithm 3, we provide a detailed explanation of the Multi-relational Prompt-based Pairwise Message Learning Mechanism. After obtaining message representations (Algorithm 3, line 3) and their respective similarity (Algorithm 3, line 4), during the training phase, we construct the pairwise cross-entropy loss based on the similarity of messages and actual event relation labels between them. Additionally, the intra-cluster loss and inter-cluster loss (Algorithm 3, line 7) are constructed by the central feature matrix (Algorithm 3, line 6) and message representations. The model is optimized through backpropagation of the total loss (Algorithm 3, line 8). During the detection phase, we update the candidate representation set and the similarity set for each message based on the obtained candidate representations of messages and their corresponding similarities. (Algorithm 3, lines 11, 12). Subsequently, we obtain the final representation of each message through the similarity-based representation aggregation approach (Algorithm 4, lines 1-3). Finally, we utilize the final representations of all messages along with their corresponding true event labels to detect events through clustering algorithms (Algorithm 4, line 5). Herein, we employ two distinct clustering methods: the distance-based K-Means (MacQueen, 1967) algorithm and the density-based HDBSCAN (Campello et al., 2013) algorithm. HDBSCAN is an enhanced version of the DBSCAN (Ester et al., 1996) algorithm, incorporating concepts from hierarchical clustering. This method does not require pre-specifying the number of clusters. Instead, it identifies the optimal clustering outcome by constructing a density hierarchy. Compared to the original DBSCAN, HDBSCAN offers greater flexibility in parameter settings, effectively adapts to data from different density regions, and exhibits stronger robustness. In practice, density-based clustering algorithms are more suitable as they do not require pre-setting the number of clusters. This allows them to better adapt to incremental data changes and meet the needs for continuous or dynamic detection.
Input: Candidate representation sets of messages: $X_{1},X_{2},X_{3},\cdots$, and their corresponding similarity sets: $Q_{1},Q_{2},Q_{3},\cdots$, ground truth labels: $L^{true}$.
Output: A set of clustered events: $e_{0},e_{1},e_{2},\cdots$.
1 for $X,Q~{}in~{}[X_{1},X_{2},X_{3},\cdots],[Q_{1},Q_{2},Q_{3},\cdots]$ do
2       $\overline{x}\leftarrow$ Eq. 7;
3       Append final message representation $\overline{x}$ into the clustering feature matrix $Matrix$;
4
5 end for
$e_{0},e_{1},e_{2},\cdots\leftarrow Cluster(Matrix,L^{true})$;
Algorithm 4 Message Representation Aggregation and Clustering.
@@ 3.6. Time Complexity
The overall time complexity of $\mathrm{RPLM}_{SED}$ is $O(N)$, where $N$ represents the number of messages. Specifically, the computational time for constructing social messages into a multi-relational graph is $O(N_{e})=O(N)$; here, $N_{e}$ is the total number of edges in the message graph. Subsequently, the time complexity of the pairwise message modeling process is $O(S\cdot N)$, where $S$ is the number of message pairs sampled for each message. Additionally, the computational time for the entire encoding process is $O(S\cdot N)$, consistent with the time complexity of message modeling. Thus, the overall time complexity of the model is $O(N_{e}+S\cdot N+S\cdot N)=O(N)$. It should be noted that, in practical scenarios, the training duration of the model is also influenced by the scale of the selected PLM.
@@ 4.1. Datasets
  • Events2012 (McMinn et al., 2013). The Events2012 dataset contains 68,841 annotated English tweets covering 503 different event categories, encompassing tweets over a consecutive 29-day period. The Event2012_100 contains 100 events with a total of 15,019 tweets, where the maximal event comprises 2,377 tweets, and the minimally has 55 tweets, with an imbalance ratio of approximately 43.
  • Events2018 (Mazoyer et al., 2020). The Events2018 includes 64,516 annotated French tweets covering 257 different event categories, with data spanning over a consecutive 23-day period. The Event2018_100 contains 100 events with a total of 19,944 tweets, where the maximal event comprises 4,189 tweets and the minimally has 27 tweets, an imbalance ratio of approximately 155.
@@
Dataset
Messages
Events
Timespan
Max. Event
Min. Event
Avg. Event
Events2012
68841
503
29 days
7133
1
137
Events2018
64516
257
23 days
12567
1
251
Events2012_100
15019
100
2377
55
150
Events2018_100
19944
100
4189
27
199
Arabic-Twitter
9070
7
2883
519
1296
4.2. Baseline
@@
  • BERT (Kenton and Toutanova, 2019): BERT is a pre-trained model for natural language processing (NLP) proposed by researchers at Google in 2018. The cornerstone of this model’s innovation lies in its utilization of the encoder architecture from the Transformer model to train context-based word embeddings, incorporating a Masked Language Model (MLM) design. Consistent with previous methodologies, we employ BERT to obtain message representations for clustering directly. BERT-FT achieves streaming event detection by applying fine-tuning to BERT. Given the inapplicability of cross-entropy in open event sets, the encoding process is optimized through online triplet loss and the implementation of proposed clustering constraints. Analogous to standard incremental event detection methodologies, BERT-FT, after its training or maintenance, detects events within message blocks inside the window.
@@ 4.3. Experimental Setting and Implementation
We implement all models using Python 3.8 and PyTorch 2.0. All experiments are conducted on a 32-core Intel Core i9-13900K@3.80GHz, with 128GB RAM and an NVIDIA GeForce RTX 4090 GPU. For PP-GCN [^1] https://github.com/RingBDStack/PPGCN, KPGNN [^2] https://github.com/RingBDStack/KPGNN, QSGNN [^3] https://github.com/RingBDStack/open-set-social-event-detection and FinEvent [^4] https://github.com/RingBDStack/FinEvent, we follow their open-source implementations. We implement EventX with Python 3.7. To ensure fairness in comparison, unless specifically noted, we uniformly apply the K-Means (MacQueen, 1967) clustering algorithm to all baselines, excluding EventX, setting the total number of categories equal to the number of event categories in the real world. Additionally, we train the FinEvent framework from scratch on the French dataset, rather than utilizing its proposed cross-lingual social message representation learning method, ‘Crlme’. Furthermore, for the DGL-DBSCAN method proposed by FinEvent, we conduct a more comprehensive comparison by further clustering the message representations generated by $\mathrm{RPLM}_{SED}$ through the HDBSCAN (Campello et al., 2013) clustering method, and conduct a comprehensive comparison and analysis of the results.
@@ 5. Results And Discussion
  • Q6: How do different components and hyperparameters in $\mathrm{RPLM}_{SED}$ influence overall effectiveness (Section 5.6, Section 5.7)?
  • Q7: How does the performance of $\mathrm{RPLM}_{SED}$ compare to the strongest baseline in a specific scenario (Section 5.8)?
@@ 5.1. Model Effectiveness
$\mathrm{RPLM}_{SED}^{\mathrm{k}}$
$\mathrm{RPLM}_{SED}^{\mathrm{d}}$
Improve
$M_{0}$ in Events 2012 NMI
.$70\pm.02$
.$68\pm.00$
.$76\pm.02$
.$79\pm.01$
.$84\pm.01$
.$\underline{.86\pm.01}$
.$86\pm.01$
.$\textbf{.87}\pm\textbf{.01}.$
.$\uparrow.01$
AMI
.$56\pm.01$
.$29\pm.00$
.$64\pm.02$
.$68\pm.01$
.$72\pm.01$
.77±.01plus-or-minus.77.01.$7$01
.$\underline{.81\pm.00}$
.$\textbf{.84}\pm\textbf{.00}.$
.$\uparrow.07$
ARI
.$20\pm.01$
.$05\pm.00$
.$34\pm.01$
.$38\pm.01$
.$40\pm.02$
.$\underline{.56\pm.02}$
.$41\pm.01$
.$\textbf{.80}\pm\textbf{.02}.$
.$\uparrow.24$
$M_{0}$ in Events 2018 NMI
.$60\pm.01$
.$57\pm.00$
.66±.03plus-or-minus.66.03.$6$03
.$71\pm.02$
.$73\pm.02$
.$78\pm.01$
.$\textbf{.93}\pm\textbf{.00}.$
.$\underline{.92\pm.00}$
.$\uparrow.15$
AMI
.$49\pm.02$
.$56\pm.00$
.$60\pm.02$
.$64\pm.02$
.$69\pm.01$
.$74\pm.01$
.$\textbf{.92}\pm\textbf{.01}.$
.$\underline{.91\pm.01}$
.$\uparrow.18$
ARI
.$17\pm.01$
.$01\pm.00$
.$27\pm.03$
.33±.02plus-or-minus.33.02.$3$02
.$35\pm.02$
.$61\pm.02$
.$\underline{.94\pm.01}$
.$\textbf{.95}\pm\textbf{.01}.$
.$\uparrow.34$

### v2

@@ 3.3.2. Multi-Head Structured Attention
(6) $q_{i,s}=\sigma(\mathbf{W}^{\mathsf{T}}\cdot(\hat{x}_{i};\hat{x}_{s})+b),$
@@ 3.3.3. Similarity-based Representation Aggregation
In the detection phase, we sample $n$ other message nodes for each message node $m_{i}$. Consequently, for each specific message $m_{i}$, we obtain a set of $n$ candidate representations $X_{i}=[\hat{x}_{i,1},\hat{x}_{i,2},\cdots,\hat{x}_{i,n}]$, along with the corresponding set of similarity $Q_{i}=[q_{i,1},q_{i,2},\cdots,q_{i,n}]$. Subsequently, select the indices from $Q_{i}$ that satisfy the conditions and average the candidate representations corresponding to these indices to obtain the final representation of $m_{i}$. We formalize this process as follows:
(7) $I_{i}=\{j\mid q_{i,j}>\alpha,j\in\{1,2,\ldots,n\}\},\quad\overline{x}_{i}=\frac{1}{|I_{i}|}\sum_{j\in I_{i}}\hat{x}_{i,j}.$
Here, $I_{i}$ represents the set of indices obtained by filtering the set $Q_{i}$, $\alpha$ is the similarity retention threshold, and $\overline{x}_{i}$ denotes the final representation of message $m_{i}$. Particularly, we choose the candidate representation corresponding to the highest similarity index in $Q_{i}$ as the final representation for those messages in the similarity set $Q_{i}$ with only one or no indices meeting the condition.
@@ 3.4.2. Clustering Constraint
Input: A social stream $S=M_{0},M_{1},M_{2},\cdots$, the ground truth label set: $L=L_{0},L_{1},L_{2},\cdots$, the window size: $w$, and the number of mini-batches $B$.
Output: Sets of events: $E_{0},E_{1},E_{2},\cdots$.
1
2for $t=0,1,2,\cdots$ do
3
4      /* Detect events from $M_{t}$/
5       if $t~{}>~{}0$ then
6             $\mathcal{G}\leftarrow$ Construct multi-relational message graph;
7             for $a=1,2,3,\cdots,n$ do
8                   $(m_{i},m_{s})\leftarrow\{m_{i},\text{Sample}(\mathcal{G}[\mathcal{M}_{t}])~{}|~{}\forall m_{i}\in M_{t}\}$;
9                   // Sampling messages that have richer connections with $m_{i}$ to form pairs.
10                   $r_{i,s}\leftarrow$ Eq. 1;
11                   $d\leftarrow$ Eq. 2;
12                   Append $d$ to $D$ (Eq. 3);
13             end for
14            $D,Train=False\Rightarrow$ Execute Algorithm 2;
15       end if
16
17      /* Initial model training or the model maintenance. */
18       if $t\%w==0$ then
19             if $t==0$ then
20                   $M,L^{gt}=M_{0},L_{0}$; // $M_{0}$ is used to train initial model $f_{0}$.
21
22            else
23                   $M=M_{t-w+1}+\cdots+M_{t}$; // Using messages in the window to maintain the model.
24                   $L^{gt}=L_{t-w+1}+\cdots+L_{t}$;
25                   $\theta_{t}^{PLM}=\zeta\cdot\theta_{t-w}^{PLM}+(1-\zeta)\cdot\theta_{0}^{PLM}$;// Parameter weighted update of PLM in $\mathrm{RPLM}_{SED}$.
27             end if
28
29            /* Pairwise message modeling. */
30             $\mathcal{G}\leftarrow$ Construct multi-relational message graph;
31             for $a=1,2,3,\cdots,y$ do
32                   $(m_{i},m^{+}_{s}),(m_{i},m^{-}_{s})\leftarrow\{m_{i},\text{Sample}(\mathcal{G}[\mathcal{M}],L^{gt})~{}|~{}\forall m_{i}\in M\}$;
33                   // Utilizing ground truth labels to choose positive and negative samples.
34                   $r_{i,s}\leftarrow$ Eq. 1;
35                   $d^{+}$, $d^{-}\leftarrow$ Eq. 2;
36                   Append $d^{+}$, $d^{-}$ to $D$ (Eq. 3);
37             end for
38            $D,Train=True\Rightarrow$ Execute Algorithm 2;
39       end if
40
41 end for
1 /* Multi-relational prompt-based pairwise message learning. */
2 for $b=1,2,3,\cdots,B$ do
3       get representations of the message and sampled message from each message pair via Eq. 4, Eq. 5;
4       Get the similarity of the two in each message pair via Eq. 6;
5       if Train then
@@ 9             Back propagate to update model parameters;
11             Update sets of each message candidate representations: $X_{1},X_{2},\cdots,X_{i}$;
12             Update corresponding similarity sets: $Q_{1},Q_{2},\cdots,Q_{i}$;
13
@@ 12             Update corresponding similarity sets: $Q_{1},Q_{2},\cdots,Q_{i}$;
17/* Similarity-based representation aggregation and Clustering. */
18 if not Train then
19       for $X,Q~{}in~{}[X_{1},X_{2},X_{3},\cdots,X_{i}],[Q_{1},Q_{2},Q_{3},\cdots,Q_{i}]$ do
20             $\overline{x}_{i}\leftarrow$ Eq. 7;
21             Append final message representation $\overline{x}_{i}$ into the clustering feature matrix $Matrix$;
22
23       end for
24      $E_{t}=\{e_{0},e_{1},e_{2},\cdots\}\leftarrow Cluster(Matrix)$;
25 end if
Algorithm 2 Model Training and Detection.
@@ 3.5. Proposed $\mathrm{RPLM}_{SED}$
As depicted in Algorithm 1, the life cycle of $\mathrm{RPLM}_{SED}$ is divided into three stages: the initial training phase, the detection phase, and the maintenance phase. During the initial training stage, we train the model using the initial message block (Algorithm 1, lines 17, 25-33). Subsequently, the model performs event detection on message blocks within the window (Algorithm 1, lines 3-13). To retain existing knowledge while adapting to new data, we perform a weighted update of the parameters in the PLMs within the model before the maintenance phase (Algorithm 1, line 21), then maintain the model by utilizing the message blocks within the window (Algorithm 1, lines 19, 20, 25-33). Once the model maintenance is completed, it will proceed to detect events in message blocks within the next window.
@@
Lines 4-11 and 25-32 of Algorithm 1 separately introduce the pairwise message modeling strategy in the training and detection stages, including message node sampling and multiple relation mapping. Lines 2-14 of Algorithm 2 explain the multi-relational prompt-based pairwise message learning mechanism. After obtaining message representations (Algorithm 2, line 3) and their respective similarity (Algorithm 2, line 4), during the training phase, we construct the pairwise cross-entropy loss based on the similarity of messages and actual event relations between them. Additionally, the intra-cluster loss and inter-cluster loss (Algorithm 2, line 7) are constructed by the central feature matrix (Algorithm 2, line 6) and message representations. The model is optimized through backpropagation of the total loss (Algorithm 2, line 8, 9). During the detection phase, we update the candidate representation set and the similarity set for each message based on the obtained candidate representations of messages and their corresponding similarities. (Algorithm 2, lines 11, 12). Subsequently, we obtain the final representation of each message through the similarity-based representation aggregation approach (Algorithm 2, lines 17-20). Finally, events in a message block are detected, utilizing all messages’ final representations through clustering algorithms (Algorithm 2, line 21). Herein, we employ two distinct clustering algorithms: the distance-based K-Means (MacQueen, 1967) and the density-based HDBSCAN (Campello et al., 2013) algorithms. HDBSCAN is an enhanced version of the DBSCAN (Ester et al., 1996) algorithm, incorporating concepts from hierarchical clustering. This method does not require pre-specifying the number of clusters. Instead, it identifies the optimal clustering outcome by constructing a density hierarchy. Compared to the original DBSCAN, HDBSCAN offers greater flexibility in parameter settings, effectively adapts to data from different density regions, and exhibits stronger robustness. In practice, density-based clustering algorithms are more suitable as they do not require pre-setting the number of clusters. This allows them to better adapt to incremental data changes and meet the needs for continuous or dynamic detection.
@@ 3.6. Time Complexity
The time complexity of $\mathrm{RPLM}_{SED}$ consists of pairwise message modeling, multi-relational prompt-based pairwise message learning mechanism, and message representation clustering. During pairwise message modeling, constructing a HIN graph and transforming it into a multi-relational message graph necessitates traversing all edges, resulting in a time complexity of $O(N_{e})$, where $N_{e}$ is the number of edges in the graph. The time complexity for sampling messages to form pairs is $O(N\cdot n)=O(N)$, where $N$ is the number of messages in the graph, and $n$ is the number of message pairs sampled for each message. Thus, the total time complexity of this phase is $O(N_{e}+N)=O(N_{e})$. In the multi-relational prompt-based pairwise message learning mechanism, firstly, embedding operations for message pair content and multi-relation sequences require time $O(N\cdot n\cdot s)=O(N)$, where $s$ is the average input sequence length. Secondly, the PLM comprises multiple layers of Transformers, where each layer’s primary source of time complexity stems from the self-attention mechanism. The time complexity in the encoding phase is $O(N\cdot n\cdot s^{2})=O(N\cdot s^{2})$. Notably, as the scale of PLM increases, higher layer counts and hidden dimensions lead to increased time complexity. Thirdly, the time complexities of the structured attention mechanism and classifier are $O(N\cdot n\cdot s)=O(N)$ and $O(N)$, respectively. Finally, the time complexity for selecting and aggregating final message representations is $O(N\cdot n)=O(N)$. Therefore, the overall time complexity of the pairwise message learning mechanism is $O(N+N\cdot s^{2}+N+N+N)=O(N\cdot s^{2})$. In the clustering stage, the time complexity for clustering message representations using the K-means algorithm is $O(I\cdot K\cdot N\cdot d)=O(N)$, where $I$ and $K$ denote the number of iterations and the given number of clusters, respectively. In contrast, the time complexity of clustering with the HDBSCAN algorithm is $O(N\cdot\log N)$.
@@ 4.1. Datasets
  • Events2012 (McMinn et al., 2013). The Events2012 dataset contains 68,841 annotated English tweets covering 503 different event categories, encompassing tweets over a consecutive 29-day period. The Events2012_100 contains 100 events with a total of 15,019 tweets, where the maximal event comprises 2,377 tweets, and the minimally has 55 tweets, with an imbalance ratio of approximately 43.
  • Events2018 (Mazoyer et al., 2020). The Events2018 includes 64,516 annotated French tweets covering 257 different event categories, with data spanning over a consecutive 23-day period. The Events2018_100 contains 100 events with a total of 19,944 tweets, where the maximal event comprises 4,189 tweets and the minimally has 27 tweets, an imbalance ratio of approximately 155.
@@
Dataset Messages Events Timespan Max. tweet count/Event Min. tweet count/Event Avg. tweet count/Event
Events2012 68841 503 29 days 7133 1 137
Events2018 64516 257 23 days 12567 1 251
Events2012_100 15019 100 - 2377 55 150
Events2018_100 19944 100 - 4189 27 199
Arabic-Twitter 9070 7 - 2883 519 1296
4.2. Baselines
@@
  • CLKD (Ren et al., 2024): CLKD is a cross-lingual social event detection framework based on knowledge distillation, jointing knowledge distillation, and GNNs to improve social event detection performance in low-resource language.
  • ETGNN (Ren et al., 2022b): ETGNN is an evidential temporal-aware graph neural network for social event detection, which integrates view-specific graph quality and temporal information uncertainty into social event detection.
  • $\mathrm{\textbf{UCL}}_{\textbf{SED}}$ (Ren et al., 2023): $\mathrm{UCL}_{SED}$ is an uncertainty-guided class imbalance learning framework for imbalanced social event detection tasks. It focuses on boundary learning in latent space and classifier learning with high-quality uncertainty estimation.
  • MVGAN (Cui et al., 2021): MVGAN is a multi-view graph attention network for social event detection. It enriches event semantics through node aggregation and multi-view fusion in heterogeneous message graphs.
@@
  • BERT (Kenton and Toutanova, 2019): BERT is a pre-trained model for natural language processing (NLP) proposed by researchers at Google in 2018. The cornerstone of this model’s innovation lies in its utilization of the encoder architecture from the Transformer model to train context-based word embeddings, incorporating a Masked Language Model (MLM) design. Consistent with previous methodologies, we employ BERT to obtain message representations for clustering directly. BERT-FT achieves streaming event detection by applying fine-tuning to BERT. Given the inapplicability of cross-entropy in open event sets, the encoding process is optimized through online triplet loss and the implementation of proposed clustering constraints. Analogous to standard incremental event detection methods, BERT-FT, after its training or maintenance, detects events within message blocks inside the window.
@@ 4.3. Experimental Setting and Implementation
We implement all models using Python 3.8 and PyTorch 2.0. All experiments are conducted on a 32-core Intel Core i9-13900K@3.80GHz, with 128GB RAM and an NVIDIA GeForce RTX 4090 GPU. For PP-GCN [^2] https://github.com/RingBDStack/PPGCN, KPGNN [^3] https://github.com/RingBDStack/KPGNN, QSGNN [^4] https://github.com/RingBDStack/open-set-social-event-detection and FinEvent [^5] https://github.com/RingBDStack/FinEvent, we follow their open-source implementations. We implement EventX with Python 3.7. To ensure fairness in comparison, unless specifically noted, we uniformly apply the K-Means (MacQueen, 1967) clustering algorithm to all baselines and $\mathrm{RPLM}_{SED}$, excluding EventX, setting the total number of categories equal to the number of event categories in the real world. Additionally, we train the FinEvent framework from scratch on the French dataset, rather than utilizing its proposed cross-lingual social message representation learning method, ‘Crlme’. Furthermore, for the DGL-DBSCAN method proposed by FinEvent, we conduct a more comprehensive comparison by further clustering the message representations generated by $\mathrm{RPLM}_{SED}$ through the HDBSCAN (Campello et al., 2013) clustering method, and conduct a comprehensive comparison and analysis of the results. Where the minimum cluster size parameter of HDBSCAN is set to 8.
@@ 5. Results And Discussion
  • Q6: How does the time consumption of $\mathrm{RPLM}_{SED}$ based on different scale PLMs (Section 5.6)?
  • Q7: How do different components and hyperparameters in $\mathrm{RPLM}_{SED}$ influence overall effectiveness (Section 5.7, Section 5.8)?
  • Q8: How does the performance of $\mathrm{RPLM}_{SED}$ compare to the strongest baseline in a specific scenario (Section 5.9)?
  • Q9: How does the performance of $\mathrm{RPLM}_{SED}$ based on different cluster algorithms (Section 5.10)?
@@ 5.1. Model Effectiveness
$M_{0}$ in Events 2012 NMI .$70\pm.02$$68\pm.00$$76\pm.02$$79\pm.01$$84\pm.01$$\underline{.86\pm.01}$$86\pm.01$$\textbf{.87}\pm\textbf{.01}.$$\uparrow.01$
AMI .$56\pm.01$$29\pm.00$$64\pm.02$$68\pm.01$$72\pm.01$$7$01 .$\underline{.81\pm.00}$$\textbf{.84}\pm\textbf{.00}.$$\uparrow.07$
ARI .$20\pm.01$$05\pm.00$$34\pm.01$$38\pm.01$$40\pm.02$$\underline{.56\pm.02}$$41\pm.01$$\textbf{.80}\pm\textbf{.02}.$$\uparrow.24$
$M_{0}$ in Events 2018 NMI .$60\pm.01$$57\pm.00$$6$03 .$71\pm.02$$73\pm.02$$78\pm.01$$\textbf{.93}\pm\textbf{.00}.$$\underline{.92\pm.00}$$\uparrow.15$
AMI .$49\pm.02$$56\pm.00$$60\pm.02$$64\pm.02$$69\pm.01$$74\pm.01$$\textbf{.92}\pm\textbf{.01}.$$\underline{.91\pm.01}$$\uparrow.18$
ARI .$17\pm.01$$01\pm.00$$27\pm.03$$3$02 .$35\pm.02$$61\pm.02$$\underline{.94\pm.01}$$\textbf{.95}\pm\textbf{.01}.$$\uparrow.34$